# Universal Efimov Scaling in the Rabi-Coupled Few-Body Spectrum

### v1

@@ Acknowledgements.
At a quantitative level, there are some differences between the results of the two models. This is primarily due to the fact that the effective range in the $R^{*}$ model modifies the two-body bound state when $a\lesssim R^{*}$, and thus (primarily) the deepest Efimov states. To check the universality in the regime where this dependence is negligible, we consider the Efimov states for $|a|\gg R^{*}$ in the $R^{*}$ model, where the two-body binding energy is essentially the same as in the $\Lambda$ model. The result is shown in Fig. S2, similar to Fig. 2(a) in the main text, where we parameterize the Efimov states by $a_{-}^{(5)}/\lambda^{5}$ rather than by $a_{-}$. We see that the difference between the two models indeed disappears for the higher excited states. However, as we approach larger drives, and start probing higher energy physics, we again see the difference between the models, as expected. In both models, the new three-body parameter asymptotes to 00 at very large drives, although the precise manner in which they asymptote is non-universal.
@@
$R^{*}$ model $8.63$ $5.57$ $5.11$ $4.94$ $4.89$ $4.87$
$\Lambda$ model $13.91$ $6.40$ $5.18$ $4.95$ $4.89$ $4.87$
@@
To gain insight into the essential physics that allows the three-body parameter to be tuned via the Rabi drive, we use the Born-Oppenheimer approximation Bethe and Peierls (1935), which is valid in the experimentally relevant case of a Cs-Li mixture where $M\gg m$. In this limit, we can first assume that the light atom is moving around two fixed heavy atoms, located at ${\bf R}_{1}$ and ${\bf R}_{2}$. The $\sigma$ component of the Rabi-coupled light atom’s wave function at position ${\bf r}$ can then be written as $\psi_{\sigma}({\bf r})$, which satisfies the non-interacting Schrödinger equation everywhere
$\begin{pmatrix}-\frac{\nabla_{\bf r}^{2}}{2m}&\frac{\Omega_{0}}{2}\\ \frac{\Omega_{0}}{2}&-\frac{\nabla_{\bf r}^{2}}{2m}+\Delta_{0}\end{pmatrix}\begin{pmatrix}\psi_{\uparrow}({\bf r})\\ \psi_{\downarrow}({\bf r})\end{pmatrix}$ $=\epsilon({\bf R}_{1},{\bf R}_{2})\begin{pmatrix}\psi_{\uparrow}({\bf r})\\ \psi_{\downarrow}({\bf r})\end{pmatrix}\,,$ (S10)
@@
$\psi_{\uparrow}({\bf r})\propto\frac{1}{|{\bf r}-{\bf R}_{i}|}-\frac{1}{a}-2mR^{*}\epsilon({\bf R}_{1},{\bf R}_{2})$ $,\qquad\text{as}\quad|{\bf r}-{\bf R}_{i}|\to 0\,.$ (S11)
Here we include the effective-range contribution to the boundary condition Levinsen and Petrov (2011). After solving Eqs. (S10) and (S11), we assume that the two heavy atoms move on the potential energy surface $\epsilon({\bf R}_{1},{\bf R}_{2})$. Due to the translational symmetry of the system, we have $\epsilon({\bf R}_{1}-{\bf R}_{2})\equiv\epsilon(R)$ with $R=|{\bf R}_{1}-{\bf R}_{2}|$ the relative separation between the two heavy particles. To define an effective potential $V_{\text{c}}(R)$ for the heavy particles, we must subtract the new scattering continuum, and thus $V_{\text{c}}(R)=\epsilon(R)-\epsilon(\infty)=\epsilon(R)-{\epsilon_{-}}$, where ${\epsilon_{-}}$ is the shifted continuum. Then, the radial Schrödinger equation for the relative motion of the two heavy particles satisfies the effective one-dimensional Schrödinger equation
@@
Figure S3: (a) Effective interaction potentials as a function of distance between the two heavy atoms, calculated within the Born-Oppenheimer approximation multiplied by $R^{2}$ to highlight the deviations from the original $-1/R^{2}$ potential. We plot the difference $\delta V(R)$ defined in Eq. (S20) at the induced resonance $01/a_{\rm{eff}}=$ 0 and $\Delta_{0}=0$ with a finite effective range $R^{*}\sqrt{m\Omega_{0}}=10^{-3}$ (pink, dotted) and zero effective range (purple, dashed). (b) Tuning the three-body parameter using Rabi coupling. We plot the energies of the first five excited Efimov states at the induced resonance with a finite effective range and $\Delta_{0}=0$ (purple dotted potential).
@@
$G^{(0)}_{\uparrow}(R)=-\frac{m}{2\pi}\frac{u^{2}e^{-{\sqrt{2m({\epsilon_{-}}-\epsilon(R))}}R}+v^{2}e^{-{\sqrt{2m({\epsilon_{+}}-\epsilon(R))}}R}}{R}\,,$ (S14)
and $f_{1}$, $f_{2}$ are constants remaining to be determined. Since the system remains unchanged under exchange of ${\bf R}_{1}$ and ${\bf R}_{2}$ we have $f_{1}=\pm f_{2}$, with the plus sign resulting in the lowest energy $\epsilon$.
@@
where we have the $T$ matrix, Eq. (5) of the main text, in the limit $M\to\infty$ where $m_{r}\to m$,
$\mathcal{T}_{\uparrow}^{-1}(E,0)=\frac{m^{2}R^{*}}{\pi}E+\frac{m}{2\pi a}-\frac{m^{3/2}}{\sqrt{2}\pi}\left(u^{2}\sqrt{{\epsilon_{-}}-E}+v^{2}\sqrt{{\epsilon_{+}}-E}\right)\,.$ (S16)
@@
$V_{0}(R)\equiv-\frac{{W}(1)^{2}}{2mR^{2}},$ (S18)
where ${W}(x)$ is the Lambert $W$ function and ${W}(1)\approx 0.567$. This attractive $1/R^{2}$ potential is the origin of the Efimovian discrete scaling symmetry Efimov (1970), and we have $\lambda=e^{\pi/s_{0}}$ with $s_{0}=\sqrt{{W}(1)^{2}M/2m-1/4}\simeq W(1)\sqrt{M/2m}$.
@@
where we have defined $\kappa=\sqrt{-2mV_{\text{c}}(R)}$ and $k_{\Omega}=\sqrt{2m\Omega_{0}}$, and we have also taken a resonant Rabi drive such that $\Delta_{0}=0$, since we are interested in exploring universal behavior with a minimal number of energy scales. Thus, we now have an equation for the potential that depends on the Rabi coupling via the parameter $k_{\Omega}R$.
@@
It is useful to quantify the difference between the Rabi coupled potential at unitarity ($01/a_{\mathrm{eff}}=$ 0) and the universal form $V_{0}(R)$ as follows
$\delta V(R)=V_{\text{c}}(R)+\frac{W(1)^{2}}{2mR^{2}}.$ (S20)
When multiplied by $R^{2}$, this resembles a scattering potential induced by the Rabi drive, as shown in Fig. S3(a). The appearance of a potential barrier at intermediate $R$ can be linked to the Rabi-induced range parameter $R^{*}_{\text{eff}}$, while the behavior at short distances is determined by the UV physics, such as the range parameter $R^{*}$. Specifically, in the case of a finite $R^{*}$, the potential departs from the universal behavior in the limit $R\to 0$, corresponding instead to $V_{\text{c}}(R)\sim\frac{1}{mRR^{*}}$, yielding a repulsive barrier in $R^{2}\delta V(R)$, as can be seen in Fig. S3(a).
@@
$-\frac{1}{M}\frac{\partial^{2}\Phi}{\partial z^{2}}+\frac{u(z)}{2m}\Phi(z)=\frac{s_{0}^{2}}{M}\Phi(z)\,,$ (S21)
where $u(z)=W(1)^{2}-e^{2z}\kappa^{2}/k_{\Omega}^{2}$. Thus, we have a 1D problem of a particle with positive energy $s_{0}^{2}/M$ experiencing an effective potential $u(z)/2m$ that vanishes in the limits $z\to\pm\infty$ (corresponding to $k_{\Omega}R\to 0$ and $k_{\Omega}R\to\infty$).
@@
$\frac{d\delta}{dz}=-\frac{M}{2ms_{0}}u(z)\sin^{2}(s_{0}z+\delta(z)).$ (S23)
@@
Figure S4: The first seven excited Efimov trimer energies at the induced resonance ($01/a_{\mathrm{eff}}=$ 0) as a function of Rabi-coupling strength, calculated with the $R^{*}$ model. Here, we have applied the scaling relation $E_{0}^{(n)}\lambda^{2n}=E_{0}$ to collapse all the binding energies to a single initial value. The light gray dashed lines represent the condition ${\epsilon_{-}}=E_{0}^{(n)}$, where the rightmost is the ground $n=0$ case. In (a) we have $\Delta_{0}=-\Omega_{0}$ and in (b) $\Delta_{0}=\Omega_{0}$.
@@
Figure S5: The coupled binding energies at the induced resonance as a ratio of the initial binding energy for various mass ratios, calculated with the $R^{*}$ model and $\Delta_{0}=0$. Here, we show the 7th state for $M/m\approx 22.1$ (blue), 9th state for $M/m\approx 44.3$ (purple), 11th state for $M/m\approx 88.6$ (pink), 13th state for $M/m\approx 177.6$ (green). The scaling factor for these mass ratios are: $4.87,3.05,2.17$ and $1.72$ respectively. Inset: The ratio of coupled to uncoupled energies calculated within the Born-Oppenheimer approximation using the variable phase approach, Eq. (S23), for the same mass ratios.

### v2

@@ Acknowledgements.
At a quantitative level, there are some differences between the results of the two models. This is primarily due to the fact that the effective range in the $R^{*}$ model modifies the two-body bound state when $a\lesssim R^{*}$, and thus (primarily) the deepest Efimov states. To check the universality in the regime where this dependence is negligible, we consider the Efimov states for $|a|\gg R^{*}$ in the $R^{*}$ model, where the two-body binding energy is essentially the same as in the $\Lambda$ model. The result is shown in Fig. S2, similar to Fig. 2(a) in the main text, where we parameterize the Efimov states by $a_{-}^{(5)}/\lambda^{5}$ rather than by $R^{*}$. We see that the difference between the two models indeed disappears for the higher excited states. However, as we approach larger drives, and start probing higher energy physics, we again see the difference between the models, as expected. In both models, the new three-body parameter asymptotes to 0 at very large drives, although the precise manner in which they asymptote is non-universal.
@@
$R^{*}$ model 8.638.638.63 5.575.575.57 5.115.115.11 4.944.944.94 4.894.894.89 4.874.874.87
$\Lambda$ model 13.9113.9113.91 6.406.406.40 5.185.185.18 4.954.954.95 4.894.894.89 4.874.874.87
@@
To gain insight into the essential physics that allows the three-body parameter to be tuned via the Rabi drive, we use the Born-Oppenheimer approximation Bethe and Peierls (1935), which is valid in the experimentally relevant case of a 133Cs -6Li mixture where $M\gg m$. In this limit, we can first assume that the light atom is moving around two fixed heavy atoms, located at ${\bf R}_{1}$ and ${\bf R}_{2}$. The $\sigma$ component of the Rabi-coupled light atom’s wave function at position ${\bf r}$ can then be written as $\psi_{\sigma}({\bf r})$, which satisfies the non-interacting Schrödinger equation everywhere
$\begin{pmatrix}-\frac{\nabla_{\bf r}^{2}}{2m_{r}}&\frac{\Omega_{0}}{2}\\ \frac{\Omega_{0}}{2}&-\frac{\nabla_{\bf r}^{2}}{2m_{r}}+\Delta_{0}\end{pmatrix}\begin{pmatrix}\psi_{\uparrow}({\bf r})\\ \psi_{\downarrow}({\bf r})\end{pmatrix}$ $=\epsilon({\bf R}_{1},{\bf R}_{2})\begin{pmatrix}\psi_{\uparrow}({\bf r})\\ \psi_{\downarrow}({\bf r})\end{pmatrix}\,,$ (S10)
@@
$\psi_{\uparrow}({\bf r})\propto\frac{1}{|{\bf r}-{\bf R}_{i}|}-\frac{1}{a}-2m_{r}R^{*}\epsilon({\bf R}_{1},{\bf R}_{2})$ $,\qquad\text{as}\quad|{\bf r}-{\bf R}_{i}|\to 0\,,$ (S11)
which includes the effective-range contribution Levinsen and Petrov (2011). Here, we have used the reduced mass $m_{r}$ rather than $m$ in order to correctly describe the two-body physics. After solving Eqs. (S10) and (S11), we assume that the two heavy atoms move on the potential energy surface $\epsilon({\bf R}_{1},{\bf R}_{2})$. Due to the translational symmetry of the system, we have $\epsilon({\bf R}_{1}-{\bf R}_{2})\equiv\epsilon(R)$ with $R=|{\bf R}_{1}-{\bf R}_{2}|$ the relative separation between the two heavy particles. To define an effective potential $V_{\text{c}}(R)$ for the heavy particles, we must subtract the new scattering continuum, and thus $V_{\text{c}}(R)=\epsilon(R)-\epsilon(\infty)=\epsilon(R)-{\epsilon_{-}}$, where ${\epsilon_{-}}$ is the shifted continuum. Then, the radial Schrödinger equation for the relative motion of the two heavy particles satisfies the effective one-dimensional Schrödinger equation
@@
Figure S3: (a) Effective interaction potentials as a function of distance between the two heavy atoms, calculated within the Born-Oppenheimer approximation multiplied by $R^{2}$ to highlight the deviations from the original $-1/R^{2}$ potential. We plot the difference $\delta V(R)$ defined in Eq. (S20) at the induced resonance $1/a_{\rm{eff}}=0$ and $\Delta_{0}=0$ with a finite effective range $R^{*}\sqrt{m\Omega_{0}}=10^{-3}$ (purple, dashed) and zero effective range (blue). (b) Tuning the three-body parameter using Rabi coupling. We plot the energies of the first five excited Efimov states at the induced resonance with a finite effective range and $\Delta_{0}=0$ (purple dashed potential in panel (a)).
@@
$G^{(0)}_{\uparrow}(R)=-\frac{m_{r}}{2\pi}\frac{u^{2}e^{-{\sqrt{2m_{r}({\epsilon_{-}}-\epsilon(R))}}R}+v^{2}e^{-{\sqrt{2m_{r}({\epsilon_{+}}-\epsilon(R))}}R}}{R}\,,$ (S14)
and $f_{1}$, $f_{2}$ are constants remaining to be determined. Since the system remains unchanged under exchange of ${\bf R}_{1}$ and ${\bf R}_{2}$ we have $f_{1}=\pm f_{2}$, with the minus sign resulting in the lowest energy $\epsilon$.
@@
where we have the $T$ matrix, Eq. (5) of the main text,
$\mathcal{T}_{\uparrow}^{-1}(E,0)=\frac{m_{r}^{2}R^{*}}{\pi}E+\frac{m_{r}}{2\pi a}-\frac{m_{r}^{3/2}}{\sqrt{2}\pi}\left(u^{2}\sqrt{{\epsilon_{-}}-E}+v^{2}\sqrt{{\epsilon_{+}}-E}\right)\,.$ (S16)
@@
$V_{0}(R)\equiv-\frac{{W}(1)^{2}}{2m_{r}R^{2}},$ (S18)
where ${W}(x)$ is the Lambert $W$ function and ${W}(1)\approx 0.567$. This attractive $1/R^{2}$ potential is the origin of the Efimovian discrete scaling symmetry Efimov (1970), and we have $\lambda=e^{\pi/s_{0}}$ with $s_{0}=\sqrt{{W}(1)^{2}M/2m_{r}-1/4}\simeq W(1)\sqrt{M/2m_{r}}$.
@@
where we have defined $\kappa=\sqrt{-2m_{r}V_{\text{c}}(R)}$ and $k_{\Omega}=\sqrt{2m_{r}\Omega_{0}}$, and we have also taken a resonant Rabi drive such that $\Delta_{0}=0$, since we are interested in exploring universal behavior with a minimal number of energy scales. Thus, we now have an equation for the potential that depends on the Rabi coupling via the parameter $k_{\Omega}R$.
@@
It is useful to quantify the difference between the Rabi-coupled potential at unitarity ($1/a_{\mathrm{eff}}=0$) and the universal form $V_{0}(R)$ as follows
$\delta V(R)=V_{\text{c}}(R)+\frac{W(1)^{2}}{2m_{r}R^{2}}.$ (S20)
When multiplied by $R^{2}$, this resembles a scattering potential induced by the Rabi drive, as shown in Fig. S3(a). The appearance of a potential barrier at intermediate $R$ can be linked to the Rabi-induced range parameter $R^{*}_{\text{eff}}$, while the behavior at short distances is determined by the UV physics, such as the underlying range parameter $R^{*}$. Specifically, in the case of a finite $R^{*}$, the potential departs from the universal behavior in the limit $R\to 0$, corresponding instead to $V_{\text{c}}(R)\sim-\frac{1}{2m_{r}RR^{*}}$, yielding a repulsive barrier in $R^{2}\delta V(R)$, as can be seen in Fig. S3(a).
@@
$-\frac{1}{M}\frac{\partial^{2}\Phi}{\partial z^{2}}+\frac{u(z)}{2m_{r}}\Phi(z)=\frac{s_{0}^{2}}{M}\Phi(z)\,,$ (S21)
where $u(z)=W(1)^{2}-e^{2z}\kappa^{2}/k_{\Omega}^{2}$. Thus, we have a 1D problem of a particle with positive energy $s_{0}^{2}/M$ experiencing an effective potential $u(z)/2m_{r}$ that vanishes in the limits $z\to\pm\infty$ (corresponding to $k_{\Omega}R\to 0$ and $k_{\Omega}R\to\infty$).
@@
$\frac{d\delta}{dz}=-\frac{M}{2m_{r}s_{0}}u(z)\sin^{2}(s_{0}z+\delta(z)).$ (S23)
@@
S1 Three-body losses
When three atoms approach each other—within distances of the order of $r_{\text{vdW}}$—they can recombine into a deeply bound dimer state. In this process, the molecular binding energy is released in the form of dimer-atom recoil kinetic energy, which is much larger than the trap depth and temperature. This process of three-body recombination is the main form of atom loss within resonant Bose gases and Bose-Fermi mixtures. We can model this loss process via an inelasticity parameter $\eta$ Braaten and Hammer (2006); Ulmanis et al. (2016a); Sykes et al. (2017), which corresponds to the addition of a small imaginary component to the phase $\delta$ and thus dictates the difference in amplitude between the incoming and outgoing waves.
To investigate the effect of Rabi coupling on the inelasticity parameter, we once again employ the variable phase approach. Specifically, we consider a complex phase given by $\delta(z)=\delta_{0}(z)+i\eta(z)$, and solve Eq. (S23). When $\eta(z)\ll\delta_{0}(z)$, Eq. (S23) can be written
$\frac{d\delta_{0}}{dz}$ $\simeq-\frac{M}{2m_{r}s_{0}}u(z)\left[\sin^{2}\left(s_{0}z+\delta_{0}(z)\right)-\eta(z)^{2}\cos\left(2s_{0}z+2\delta_{0}(z)\right)\right]\,,$ (S24a)
$\frac{d\eta}{dz}$ $\simeq-\frac{M}{2m_{r}s_{0}}u(z)\eta(z)\sin\left(2s_{0}z+2\delta_{0}(z)\right).$ (S24b)
In the limit $z\to-\infty$, we have the modified short-distance boundary condition involving the underlying inelasticity parameter: $\delta(z)\to\delta_{-}+i\eta$. Then, the phase to the right of the potential becomes $\delta(z)\to\delta_{+}+i\eta_{\Omega}$, where $\eta_{\Omega}$ gives a measure of the Rabi-dressed loss rate, while the ratio of trimer energies is once again $e^{2(\delta_{+}-\delta_{-})/s_{0}}$. From Eq. (S24a), we can see that the presence of the inelasticity parameter does not affect the log-periodic dependence on the Rabi drive, but it can change the amplitude of the oscillations in the trimer energies.
Figure S4: Effect of three-body losses on the Rabi-dressed three-body parameter for the bare inelasticity $\eta\to 0^{+}$ (black, solid), $\eta=0.3$ (green, dotted), $\eta=0.6$ (blue, dashed) and $\eta=0.9$ (pink, dot-dashed). For typical experimental parameters Ulmanis et al. (2016b), this corresponds to $\eta\approx 0.6$ or $\eta\approx 0.9$. (a) The effect on the ratio of a given trimer’s binding energy as a function of the Rabi coupling. (b) The Rabi-dressed inelasticity parameter as a ratio of the bare inelasticity parameter.
In Fig. S4(a), we show the effect of the inelasticity parameter on the ratio of the binding energy of a given trimer in the universal regime ($|\delta E|\ll\Omega_{0}$), with and without Rabi coupling, for various values of $\eta$. We find that while the average value of the ratio does not change, the amplitude of the log-periodic oscillations increases with larger inelasticity. Likewise, Fig. S4(b) shows the dressed inelasticity parameter, $\eta_{\Omega}$, which is similarly log-periodic with increasing Rabi coupling, with a mean value of the bare parameter. Therefore, we expect that the Rabi drive would have a minimal effect on the underlying collisional losses.
S2 Intraspecies Interactions
For heteronuclear systems, the three-body parameter is not only dependent on the details of the pairwise interactions and the mass ratio, but can also depend on the intraspecies interactions Häfner et al. (2017). It is therefore interesting to investigate the manner in which such interactions can modify our results.
To this end, as in Ref. Ulmanis et al. (2016a); Häfner et al. (2017), we model the intraspecies interactions via a hard-core van der Waals potential,
$V_{\mathrm{BB}}(R)=\begin{cases}\infty&R\leq R_{0}\\ -C_{6}/R^{6}&R>R_{0}\end{cases}\,,$ (S25)
where $C_{6}$ is the van der Waals coefficient. Without the inclusion of the attractive van der Waals potential, the Cs-Cs scattering length is simply $a_{\mathrm{BB}}=R_{0}$. In general, $a_{\mathrm{BB}}$ is related to $R_{0}$ and $C_{6}$ via the equation Gribakin and Flambaum (1993); Tran et al. (2021)
$\frac{Y_{1/4}(2R_{\mathrm{vdW}}/R^{2}_{0})}{J_{1/4}(2R_{\mathrm{vdW}}/R^{2}_{0})}=1-\sqrt{2}\frac{a_{\mathrm{BB}}}{r_{\mathrm{vdW}}}\frac{\Gamma(5/4)}{\Gamma(3/4)},$ (S26)
where $J_{\alpha}(x)$ [$Y_{\alpha}(x)$] are the Bessel functions of the first [second] kind, $\Gamma(x)$ is the Gamma function and $R_{\mathrm{vdW}}=1/2(MC_{6})^{1/4}$ is the van der Waals length. Equation (S26) has multiple solutions for a given $a_{\mathrm{BB}}$, corresponding to different numbers of bound states in the attractive potential. Therefore, the specific value of $R_{0}$ not only sets $a_{\mathrm{BB}}$, but also the total number of supported bound states. Following the work of Refs. Ulmanis et al. (2016a); Häfner et al. (2017), we choose $R_{0}$ such that the potential supports two Cs dimer states.
In the absence of Rabi coupling, it is known that for small positive $a_{\text{BB}}$ and negative heavy-light interactions, the deepest bound Efimov trimers do not enter the three-body continuum Häfner et al. (2017); Ulmanis et al. (2016a). Instead, they dissociate into the light atom plus heavy-heavy dimer continuum, causing the three-body parameter to be drastically shifted from the models without Cs-Cs interaction. Therefore, we now include the intraspecies interactions in the Rabi-coupled CsLi three-body system using the effective potential $V_{\text{c}}$ and the intraspecies interactions in Eq. (S26).
Figure S5: Energies of the first five excited Efimov states at the induced resonance for $\Delta_{0}=0$. Here, we have included intraspecies (Cs-Cs) scattering, with a scattering length $a_{\mathrm{BB}}=R^{*}/2$. For typical experimental parameters Tung et al. (2014), this corresponds to $a_{\mathrm{BB}}\approx 100a_{0}$.
Within the Born-Oppenheimer approximation, our effective BB Schrödinger equation is modified to be:
$\left[-\frac{1}{M}\frac{\partial^{2}}{\partial R^{2}}+V_{\text{c}}(R)+V_{\text{BB}}(R)\right]\Psi(R)=\delta E\Psi(R)\,.$ (S27)
Solving Eq. (S27) allows us to capture the interplay between the three relevant scales, namely $\Omega_{0}$, $R^{*}$ and $a_{\text{BB}}$. In Fig. S5 we perform the same analysis as in Fig. 2(a) of the main text and Fig. S3(b) above, where now we look at the behavior of the Rabi-dressed energies as a function of increasing drive strength for the specific value $a_{\text{BB}}=R^{*}/2$. As expected, the ground state is now so deeply bound that it is completely insensitive to the drive. On the other hand, evidently the excited Efimov states look qualitatively very similar to those in Fig. S3(b), albeit with a different energy in the limit $\Omega_{0}\to 0$, indicating that (as expected) the system has a modified three-body parameter. Crucially, the relative energy shifts due to Rabi coupling are essentially exactly the same as in the case where $a_{\mathrm{BB}}=0$, indicating that the Rabi drive modifies the three-body parameter in precisely the same manner. Therefore, we see that no matter what sets the three-body parameter, the Rabi-dressing universally shifts the original three-body parameter. This justifies the omission of $a_{\mathrm{BB}}$ from the results presented in the main text.
@@
Figure S6: The first seven excited Efimov trimer energies at the induced resonance ($1/a_{\mathrm{eff}}=0$) as a function of Rabi-coupling strength, calculated with the $R^{*}$ model. Here, we have applied the scaling relation $E_{0}^{(n)}\lambda^{2n}=E_{0}$ to collapse all the binding energies to a single initial value. The light gray dashed lines represent the condition ${\epsilon_{-}}=E_{0}^{(n)}$, where the rightmost is the ground $n=0$ case. In (a) we have $\Delta_{0}=-\Omega_{0}$ and in (b) $\Delta_{0}=\Omega_{0}$.
@@
Figure S7: The coupled binding energies at the induced resonance as a ratio of the initial binding energy for various mass ratios, calculated with the $R^{*}$ model and $\Delta_{0}=0$. Here, we show the 7th state for $M/m\approx 22.1$ (blue), 9th state for $M/m\approx 44.3$ (purple), 11th state for $M/m\approx 88.6$ (pink), 13th state for $M/m\approx 177.6$ (green). The scaling factor for these mass ratios are: 4.87,3.05,2.174.87,3.05,2.174.87 , 3.05 , 2.17 and 1.721.721.72 respectively. Inset: The ratio of coupled to uncoupled energies calculated within the Born-Oppenheimer approximation using the variable phase approach, Eq. (S23), for the same mass ratios.